# Restricted Path Characteristic Function Determines the Law of Stochastic Processes

### v2

@@ 6.2. Numerical experiments on hypothesis testing
We apply the trainable empirical RPCFD in (38) to the task of hypothesis testing on laws of stochastic processes. The empirical RPCFD acts as the statistic. With an abuse of notation, suppose $X$ and $Y$ are two stochastic processes in the Euclidean space. The hypotheses of our two-sample permutation test are:
@@
where $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$ denote the laws of $X$ and $Y$, respectively.
Denoted by $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{m}\}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}$ two sets of i.i.d. samples drawn respectively from $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$. Given the statistic $T(\mathbf{X},\mathbf{Y})$ (i.e., the empirical RPCFD), we apply all possible the permutations $\sigma\in\Sigma_{m+n}$ to $\mathbf{Z}:=(\mathbf{X},\mathbf{Y})$ and obtain the set $\mathcal{T}$ via
$\mathcal{T}:=\bigg{\{}T(\mathbf{Z}_{\sigma(1):\sigma(m)},\mathbf{Z}_{\sigma(m+1):\sigma(m+n)})\;\left|\;\sigma\in\Sigma_{m+n}\bigg{\}}\right..$
Under $H_{0}$, one can uniformly sample from $\mathcal{T}$ to approximate the distribution of $T(\mathbb{P}_{X},\mathbb{P}_{Y})$. Then, given the significance level $\alpha$, the null hypothesis $H_{0}$ will be rejected if $T(\mathbf{X},\mathbf{Y})$ exceeds the $(1-\alpha)\%$ quantile of $\mathcal{T}$, otherwise accepted. More details on the two-sample permutation test can be found in e.g. [26, §7].
In practice, we implicitly assume that the conditions of Theorem 5.3 are satisfied for $X$ and $Y$. This assumption is validated by the fact that in practice, only discrete data points are available and interpolation methods have to be applied. Keeping this in mind, we follow the workflow of [15]. Specifically, given the training data, which is i.i.d. generated under $H_{1}$, we first optimise the parameters of the empirical RPCFD via (38). Then, we use the optimised empirical RPCFD to conduct the two-sample permutation test on another out-of-sample test dataset, whose samples are generated under either $H_{0}$ or $H_{1}$. Here we use the test power (true rejection, the higher the better) and Type-I error (false rejection, the lower the better) as the metrics of the performance of our optimised empirical RPCFD.
Following [15, Example B.12], we set $X:=\tilde{B}_{[0,T]}$ to be the time-augmented 3-dimensional Brownian Motion (BM), and $Y:=\tilde{B}^{h}_{[0,T]}$ to be the time-augmented 3-dimensional fractional Brownian Motion (fBM) with the Hurst parameter $h\in]0,1[$. Note that $\mathbb{P}_{X}\neq\mathbb{P}_{Y}$ when $h\neq 0.5$.
To benchmark, we consider other distance-based statistics (i.e., PCFD), as well as signature-based MMDs [26]. Detailed descriptions of the experiment hyperparameters and results can be found in Appendices A and B. For readers interested in reproducing the numerical experiments, we have made the code implementation publicly available at the GitHub repository: https://github.com/DeepIntoStreams/UnitaryDevInversion.
RPCFD vs. PCFD
We compare our proposed trainable empirical RPCFD with PCFD proposed in [15]. Recall $K$ and $k$ from §6.1. In the experiments, we set $K=8$ and $k=5$. As shown in Figure 6.1 (Upper Left), the overall test power curves against various $h$ for RPCFD and PCFD are close. In the two cases $h=0.45$ and $h=0.55$, RPCFD achieves an obviously superior test power over PCFD, indicating that restricting the linear maps to $\mathfrak{t}(k,\mathbb{R})$ effectively narrows the search space, thereby making the learning more efficient. Moreover, due to the sparsity of the linear maps in RPCFD (and that RPCFD only uses real numbers), the training time of RPCFD is also notably reduced by $22.9\%$ on average; see the upper right panel of Figure 6.1.
Figure 6.1. (Upper Left) Test power of trained RPCFD and PCFD, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I error of trained RPCFD and PCFD. (Bottom) Training time of RPCFD and PCFD over 500 iterations. Here fix $K=8$ and $k=5$.
The sensitivity analysis of RPCFD when $|h-0.5|\in[0.01,0.04]$ is plotted in Figure 6.2 below. In general, increasing the matrix order $k$ improves the discriminative power of RPCFD. However, the larger $k$ may have the potential issues of overfitting and high computational cost. Therefore, in practice, one should choose the matrix order $k$ by hyperparameter tuning.
Figure 6.2. Sensitivity analysis of RPCFD (Left: mean test power over 5 runs; Right: mean Type-I error over 5 runs). The $y$-axis and $x$-axis represent the power of the trained RPCFD and the Hurst parameter $h$, respectively. Here fix $K=8$.
Comparison with Sig-MMDs
In Appendix B (Tables B.1 and B.2), the development-based distances RPCFD, OPCFD and PCFD outperform the signature-based MMDs consistently for varying $h$ in terms of the test power. In particular, when $h\in\{0.45,0.55\}$, the development-based distances improve the test power from $[0.15,0.17]$ to $[0.95,0.99]$. Moreover, the inference time of Sig-MMD-based methods scales quadratically with the sample size, while the dependency is linear for development-based methods; see Table B.3 for the inference time of each method across different sample sizes. For example, when sample size $m=n=200$, the inference time of signature MMDs is over 10 times longer than that of the distance-based methods.
7. Conclusion
This paper provides a new, constructive proof for solving the generalised moment problem on tensor algebra. On the theoretical side, it enables us to generalise the classical results on the unique determination of signature by unitary development [7] — both by relaxing the condition on infinite radius of convergence (ROC) to nonzero ROC, and by extending to Cartan developments into a variety of matrix Lie groups. This motivates us to propose a novel, efficient metric for probability measures on the path space, which features on the sparsity of the linear maps involved and hence brings about computational benefits in dimension reduction. We demonstrate its practical power in hypothesis testing for the law of stochastic processes. It would be worthwhile to investigate in the future its applications to synthetic time series generation.
Appendices
Appendix A Numerical Experiment Details
Permutation test
We first provide a detailed description of the algorithm for estimating test power and Type-I error in the two-sample permutation test in §6.
$\alpha\in(0,1)$: significance level; $N>0$: # of experiments; $M>0$: # of permutations; $\mathbf{X}$: samples from distribution $\mu$; $\mathbf{Y}$: samples from distribution $\nu$; $m>0$: sample size of $\mathbf{X}$; $n>0$: sample size of $\mathbf{Y}$; $T$: test statistic function; $H_{0}\in\{1,0\}$: whether the null hypothesis is true or false ($H_{0}=1$ if $\mu=\nu$, otherwise $H_{0}=0$)
$\mathbf{Z}\leftarrow$ Concatenate$(\mathbf{X},\mathbf{Y})$
@@ Appendix A Numerical Experiment Details
$i\leftarrow 1$
while $i\leq N$ do
     $\mathcal{T}\leftarrow$ EmptyList
     $j\leftarrow 1$
     while $j\leq M$ do
         $\sigma\sim$ Permutation($\{1,2,\cdots,m+n\}$)
         $T_{\sigma}\leftarrow T(\{\mathbf{Z}_{\sigma(1)},\mathbf{Z}_{\sigma(2)},\cdots,\mathbf{Z}_{\sigma(m)}\},\{\mathbf{Z}_{\sigma(m+1)},\cdots,\mathbf{Z}_{\sigma(m+n)}\})$
         $\mathcal{T}\text{.append}(T_{\sigma})$
         $j\leftarrow j+1$
     end while
     if $T(\mathbf{X},\mathbf{Y})>(1-\alpha)\%$ quantile of $\mathcal{T}$ then
@@
     end if
     $i\leftarrow i+1$
end while
@@
if $H_{0}$ then
[…]
     return Type_I_error
else
[…]
     return test_power
end if
Algorithm 1 Estimating the test power/Type-I error of the permutation test
Implementation details
We next provide the full details of implementation of the numerical experiments in §6, including the choice of hyperparameters and the training procedure.
@@
For the signature MMDs, we consider the linear and RBF kernels (see [26, §7.1] for more details). For the RBF kernel, the fine-tuned $2\sigma^{2}$ is set to be 0.1.
@@ Appendix B Supplementary numerical results
As shown in Figure B.1, OPCFD has a close performance as PCFD while achieving an average time reduction of about 31.0%.
Figure B.1. (Upper Left) Test power of trained OPCFD and PCFD, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I error of trained OPCFD and PCFD. (Bottom) Training time of OPCFD and PCFD over 500 iterations. Here fix $K=8$ and $k=5$.
@@ 0.25 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$
0.45 $\textbf{0.99}\pm 0.02$ $0.92\pm 0.07$ $0.74\pm 0.17$ $0.04\pm 0.04$ $0.15\pm 0.05$
0.475 $0.06\pm 0.02$ $0.12\pm 0.09$ $\textbf{0.14}\pm 0.05$ $0.01\pm 0.02$ $0.04\pm 0.02$
0.525 $0.08\pm 0.02$ $0.1\pm 0.04$ $\textbf{0.16}\pm 0.05$ $0.05\pm 0.02$ $0.07\pm 0.04$
@@
0.65 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $0.14\pm 0.05$ $\textbf{1}\pm 0$
@@
Table B.1. Test power of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1.
@@
0.2 $0.07\pm 0.05$ $0.08\pm 0.07$ $0.04\pm 0.05$ $0.04\pm 0.09$ $0.04\pm 0.04$
0.25 $0.03\pm 0.02$ $0.02\pm 0.04$ $0.05\pm 0.05$ $0.03\pm 0.03$ $0.04\pm 0.04$
0.3 $0.05\pm 0$ $0.02\pm 0.02$ $0.04\pm 0.04$ $0.04\pm 0.04$ $0.01\pm 0.02$
0.35 $0.04\pm 0.04$ $0.05\pm 0.03$ $0.01\pm 0.02$ $0.03\pm 0.04$ $0.07\pm 0.07$
0.4 $0.04\pm 0.04$ $0.04\pm 0.04$ $0.04\pm 0.02$ $0.06\pm 0.05$ $0.04\pm 0.04$
0.425 $0.08\pm 0.07$ $0.05\pm 0.04$ $0.01\pm 0.02$ $0.03\pm 0.03$ $0.04\pm 0.04$
0.45 $0.08\pm 0.02$ $0.08\pm 0.09$ $0.09\pm 0.07$ $0.05\pm 0.04$ $0.02\pm 0.03$
0.475 $0.02\pm 0.04$ $0.07\pm 0.07$ $0.05\pm 0.03$ $0.05\pm 0.04$ $0.07\pm 0.06$
0.525 $0.09\pm 0.07$ $0.02\pm 0.02$ $0.04\pm 0.06$ $0.03\pm 0.03$ $0.04\pm 0.02$
0.55 $0.03\pm 0.04$ $0.05\pm 0.08$ $0.06\pm 0.05$ $0.07\pm 0.06$ $0.05\pm 0.05$
0.575 $0.04\pm 0.04$ $0.04\pm 0.04$ $0.1\pm 0.03$ $0.02\pm 0.03$ $0.06\pm 0.07$
0.6 $0.06\pm 0.04$ $0.04\pm 0.04$ $0.04\pm 0.04$ $0.05\pm 0.04$ $0.05\pm 0.04$
0.65 $0.02\pm 0.02$ $0.07\pm 0.02$ $0.1\pm 0.07$ $0.07\pm 0.06$ $0.03\pm 0.04$
0.7 $0.08\pm 0.05$ $0.04\pm 0.04$ $0.06\pm 0.05$ $0.05\pm 0.05$ $0.03\pm 0.04$
0.75 $0.05\pm 0.03$ $0.07\pm 0.07$ $0.05\pm 0.04$ $0.06\pm 0.07$ $0.04\pm 0.07$
0.8 $0.01\pm 0.02$ $0.05\pm 0.04$ $0.06\pm 0.02$ $0.07\pm 0.07$ $0.04\pm 0.04$
Table B.2. Type-I error of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1.
@@
1024 $92.46\pm 4.93$ $86.53\pm 3.07$ $123.71\pm 1.56$- --
Table B.3. Inference time of the permutation test across different sample sizes ($m=n$) and the training time of RPCFD, OPCFD and PCFD before conducting the permutation test. The result is in the form of mean $\pm$ std over $5$ runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $0.12\sigma^{2}=$ 0.1. Here fix $h=0.45$.
Acknowledgement. We would like to thank Dr. Lei Jiang for his help with running the codes on the servers. SL and HN are deeply indebted to Professor Haojie Ren for insightful discussions on hypothesis testing.
Funding sources. The research of SL is supported by NSFC Projects 12201399, 12331008, and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, the National Key Research &\&& Development Program 2023YFA1010900, and the Shanghai Frontier Research Institute for Modern Analysis. HN is supported by the EPSRC under the program grant EP/S026347/1 and the Alan Turing Institute under the EPSRC grant EP/N510129/1. Both SL and HN are partially funded by the SJTU-UCL joint seed fund WH610160507/067 and the Royal Society International Exchanges 2023 grant (IEC/NSFC/233077).

### v3

@@ 6.2. Numerical experiments on hypothesis testing
We apply the trainable empirical RPCFD in (38) to the task of hypothesis testing on distributions of stochastic processes. More specifically, we first train the empirical RPCFD (38) on the training set. We then use the trained empirical RPCFD as the statistic to conduct a two-sample permutation test on the test set. Finally, we use the test power (the higher the better) and Type-I error (the lower the better) as the performance metrics.
Two-sample permutation test
Suppose $X$ and $Y$ are two stochastic processes with distributions $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$, respectively. The hypotheses of the two-sample permutation test are:
$H_{0}:\mathbb{P}_{X}=\mathbb{P}_{Y}\quad\text{(null hypothesis)},$
$H_{1}:\mathbb{P}_{X}\neq\mathbb{P}_{Y}\quad\text{(alternative hypothesis)}.$
The test power is defined as the probability of correctly rejecting $H_{0}$ when $H_{1}$ is true, and the Type-I error denotes the probability of falsely rejecting $H_{0}$ when $H_{0}$ is true.
Denote by $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{m}\}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}$ two sets of i.i.d. samples drawn respectively from $\mathbb{P}_{X}$ and $\mathbb{P}_{Y}$. Given the statistic $T(\mathbf{X},\mathbf{Y})$ (here $T$ is the trained empirical RPCFD), we apply all possible permutations $\sigma\in\Sigma_{m+n}$ to $\mathbf{Z}:=(\mathbf{X},\mathbf{Y})$ and obtain
$\mathcal{T}:=\bigg{\{}T\big{(}\mathbf{Z}_{\sigma(1):\sigma(m)},\mathbf{Z}_{\sigma(m+1):\sigma(m+n)}\big{)}:\;\sigma\in\Sigma_{m+n}\bigg{\}}.$
Under $H_{0}$, one can uniformly sample from $\mathcal{T}$ to approximate the distribution of $T(\mathbb{P}_{X},\mathbb{P}_{Y})$. Given the significance level $\alpha$, the null hypothesis $H_{0}$ will be rejected if $T(\mathbf{X},\mathbf{Y})$ exceeds the $(1-\alpha)\%$ quantile of $\mathcal{T}$, and otherwise accepted.
The algorithm of our experiments is in Appendix A. See e.g., Chevyrev–Oberhauser [28, §7] for more details on two-sample permutation tests.
Constructing training and test sets
Following [16, Example B.12], we set $X:=\tilde{B}_{[0,T]}$ to be the time-augmented 3-dimensional Brownian Motion (BM), and $Y:=\tilde{B}^{h}_{[0,T]}$ to be the time-augmented 3-dimensional fractional Brownian Motion (fBM) with the Hurst parameter $h\in]0,1[$. Note that $\mathbb{P}_{X}\neq\mathbb{P}_{Y}$ when $h\neq 0.5$. The training data are i.i.d. generated under $H_{1}$, and the test data (i.e., out-of-sample data) are i.i.d. generated under either $H_{0}$ or $H_{1}$, depending on the choice of performance metrics. See Appendix A for settings of hyperparameters.
Remark 6.4.
For any $\varepsilon>0$, BM is of finite $(2+\epsilon)$-variation and fBM is of finite $(1/h+\epsilon)$-variation. Neither is of finite 2-variation when $h\leq 0.5$. Our numerics indicate that the trained empirical EPCFD is a powerful statistic to separate different measures on the highly oscillatory paths.
RPCFD vs. PCFD
We first compare the performance of our proposed trainable empirical RPCFD (38) with PCFD proposed in [16]. In our experiments, we set $K=8$ and $k=5$, and both distances are trained on the same training set by exactly the same method. As shown in Figure 6.1 (Upper Left), the overall test power curves against various $h$ of RPCFD and PCFD are close. For $h=0.45$ and $h=0.55$, RPCFD achieves an obviously superior test power over PCFD. Figure 6.1 (Upper Right) indicates that the two distances have comparable performances in Type-I error, while the bottom panel shows that the training time of RPCFD has been notably reduced ($\approx 22.9\%$ on average) compared with that of PCFD.
Remark 6.5.
The improved performance of RPCFD over PCFD indicates that restricting the linear maps to $\mathfrak{t}(k,\mathbb{R})$ narrows down the search space and hence makes the learning more efficient. On the other hand, the training time of RPCFD is also reduced, since $\mathfrak{t}(k,\mathbb{R})$ has only real entries and thus simplifies the computation (in contrast, all parameters in PCFD are complex) by a factor of $2$. In Figure B.1, we repeat the experiments with RPCFD replaced by OPCFD (cf. Remark 6.2) and report the experiment results. Similar results are observed, which supports our analysis.
The sensitivity analysis of RPCFD when $|h-0.5|\in[0.01,0.04]$ is plotted in Figure 6.2. In general, increasing the matrix order $k$ improves the discriminative power. However, a larger value of $k$ may have the potential issues of overfitting and higher computational costs. Therefore, in practice, one should choose $k$ by hyperparameter tuning on a cross-validation set.
Comparison with Sig-MMDs
In Appendix B, we provide a comprehensive summary of the test powers, Type-I errors, and time complexities of all the distances/models (i.e., RPCFD, OPCFD, PCFD and signature-based MMDs) in Tables B.1, B.2, and B.3. According to Table B.1, the development-based distances (i.e., RPCFD, OPCFD, and PCFD) outperform the signature-based MMDs consistently for various $h$ in test power. In particular, when $h\in\{0.45,0.55\}$, the development-based distances improve the test power from $[0.15,0.17]$ to $[0.95,0.99]$. Table B.2 shows that the distances/models have close performances in Type-I error. In Table B.3, we observe that the inference time of signature-based MMDs scales quadratically with respect to the sample size, while the dependency is linear for development-based methods (i.e., RPCFD, OPCFD, and PCFD). For example, when sample size $m=n=200$, the inference time of signature-based MMDs is over 10 times longer than that of the distance-based methods.
Figure 6.1. (Upper Left) Test powers of the trained RPCFD and PCFD against various $h$, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I errors of the trained RPCFD and PCFD. (Bottom) Training time of RPCFD and PCFD over 500 iterations. We fix $K=8$ and $k=5$ in all experiments.
Figure 6.2. Sensitivity analysis of RPCFD (Left: mean test power over 5 runs; Right: mean Type-I error over 5 runs). The $y$-axis and $x$-axis represent the power of the trained RPCFD and the Hurst parameter $h$, respectively. We fix $K=8$ in all experiments.
Acknowledgement. SL and HN are deeply indebted to Prof. Haojie Ren for insightful discussions on hypothesis testing. All the authors thank Dr. Lei Jiang for his help with running the codes on servers.
Funding sources. The research of SL is supported by NSFC Projects 12201399, 12331008, and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, the National Key Research &\& Development Programs 2023YFA1010900 and 2024YFA1014900, Shanghai Rising-Star Program 24QA2703600, and the Shanghai Frontier Research Institute for Modern Analysis. HN is supported by the EPSRC under the program grant EP/S026347/1 and the Alan Turing Institute under the EPSRC grant EP/N510129/1. Both SL and HN are partially funded by the SJTU-UCL joint seed fund WH610160507/067 and the Royal Society International Exchanges 2023 grant (IEC/NSFC/233077).
Appendix A Supplementary details of numerical experiments
A.1. Algorithm of two-sample permutation test in §6.2
@@ Appendix A Supplementary details of numerical experiments
5:while $i\leq N$ do
6:  $\mathcal{T}\leftarrow$ EmptyList
7:  $j\leftarrow 1$
8:  while $j\leq M$ do
9:   $\sigma\sim$ Permutation($\{1,2,\cdots,m+n\}$)
10:   $T_{\sigma}\leftarrow T(\{\mathbf{Z}_{\sigma(1)},\mathbf{Z}_{\sigma(2)},\cdots,\mathbf{Z}_{\sigma(m)}\},\{\mathbf{Z}_{\sigma(m+1)},\cdots,\mathbf{Z}_{\sigma(m+n)}\})$
11:   $\mathcal{T}\text{.append}(T_{\sigma})$
12:   $j\leftarrow j+1$
13:  end while
14:  if $T(\mathbf{X},\mathbf{Y})>(1-\alpha)\%$ quantile of $\mathcal{T}$ then
15:   $\text{num\_rejections}\leftarrow\text{num\_rejections}+1$
16:  end if
17:  $i\leftarrow i+1$
18:end while
@@
20:if $H_{0}$ then
21:  $\text{Type\_I\_error}\leftarrow\text{ratio}$
22:  return Type_I_error
23:else
24:  $\text{test\_power}\leftarrow\text{ratio}$
25:  return test_power
26:end if
Algorithm 1 Estimate the test power/Type-I error of the permutation test
A.2. Hyperparameters and training procedure in §6.2
@@
For the signature MMDs, we consider the linear and RBF kernels (see [28, §7.1] for more details). For the RBF kernel, the fine-tuned $2\sigma^{2}$ is set to be 0.1.
@@ Appendix B Supplementary numerical results
As shown in Figure B.1, OPCFD shows a comparable performance to PCFD while achieving an average time reduction of $\approx 31.0\%$.
Figure B.1. (Upper Left) Test powers of the trained OPCFD and PCFD against various $h$, where the shaded area represents mean $\pm$ std over 5 runs. (Upper Right) Type-I errors of the trained OPCFD and PCFD. (Bottom) Training time of OPCFD and PCFD over 500 iterations (time reduction of OPCFD $\approx 31.0\%$ on average). We fix $K=8$ and $k=5$ in all experiments.
@@ 0.25 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$
0.45 $\textbf{0.99}\pm 0.02$ $0.92\pm 0.07$ $74\pm 0.17 0.0$$04\pm 0.04$ $0.15\pm 0.05$
0.475 $0.06\pm 0.02$ $0.12\pm 0.09$ $\textbf{0.14}\pm 0.05$ $01\pm 0.02 0.0$$04\pm 0.02$
0.525 $0.08\pm 0.02$ $0.1\pm 0.04$ $\textbf{0.16}\pm 0.05$ $05\pm 0.02 0.0$$07\pm 0.04$
@@
0.65 $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $\textbf{1}\pm 0$ $4\pm 0.05 1±0\textbf{1}\pm 0$
@@
Table B.1. Test power of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $2\sigma^{2}=0.1$.
@@
0.2 $07\pm 0.05 0.0$$8\pm 0.07 0.0$$4\pm 0.05 0.0$$4\pm 0.09 0.0$$04\pm 0.04$
0.25 $03\pm 0.02 0.0$$2\pm 0.04 0.0$$5\pm 0.05 0.0$$3\pm 0.03 0.0$$04\pm 0.04$
0.3 $05\pm 0 0.0$$2\pm 0.02 0.0$$4\pm 0.04 0.0$$4\pm 0.04 0.0$$01\pm 0.02$
0.35 $04\pm 0.04 0.0$$5\pm 0.03 0.0$$1\pm 0.02 0.0$$3\pm 0.04 0.0$$07\pm 0.07$
0.4 $\pm 0.04 0.04±0.04$$\pm 0.04$ $04\pm 0.02 0.0$$6\pm 0.05 0.0$$04\pm 0.04$
0.425 $08\pm 0.07 0.0$$5\pm 0.04 0.0$$1\pm 0.02 0.0$$3\pm 0.03 0.0$$04\pm 0.04$
0.45 $08\pm 0.02 0.0$$8\pm 0.09 0.0$$9\pm 0.07 0.0$$5\pm 0.04 0.0$$02\pm 0.03$
0.475 $02\pm 0.04 0.0$$7\pm 0.07 0.0$$5\pm 0.03 0.0$$5\pm 0.04 0.0$$07\pm 0.06$
0.525 $09\pm 0.07 0.0$$2\pm 0.02 0.0$$4\pm 0.06 0.0$$3\pm 0.03 0.0$$04\pm 0.02$
0.55 $03\pm 0.04 0.0$$5\pm 0.08 0.0$$6\pm 0.05 0.0$$7\pm 0.06 0.0$$05\pm 0.05$
0.575 $\pm 0.04 0.04±0.04$$\pm 0.04$ $1\pm 0.03 0.0$$2\pm 0.03 0.0$$06\pm 0.07$
0.6 $06\pm 0.04 0.0$$4\pm 0.04 0.0$$4\pm 0.04 0.0$$5\pm 0.04 0.0$$05\pm 0.04$
0.65 $02\pm 0.02 0.0$$07\pm 0.02$ $1\pm 0.07 0.0$$7\pm 0.06 0.0$$03\pm 0.04$
0.7 $08\pm 0.05 0.0$$4\pm 0.04 0.0$$6\pm 0.05 0.0$$5\pm 0.05 0.0$$03\pm 0.04$
0.75 $05\pm 0.03 0.0$$7\pm 0.07 0.0$$5\pm 0.04 0.0$$6\pm 0.07 0.0$$04\pm 0.07$
0.8 $01\pm 0.02 0.0$$5\pm 0.04 0.0$$6\pm 0.02 0.0$$7\pm 0.07 0.0$$04\pm 0.04$
Table B.2. Type-I error of the distances when $h\neq 0.5$ in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $2\sigma^{2}=0.1$.
@@
1024 $92.46\pm 4.93$ $86.53\pm 3.07$ $123.71\pm 1.56$ - -
Table B.3. Inference time of the permutation test across different sample sizes and the training time of RPCFD, OPCFD and PCFD before conducting the permutation test. The result is in the form of mean $\pm$ std over 5 runs. For RPCFD, OPCFD and PCFD, fix $K=8$ and $k=5$. For the RBF signature MMD, fix $2\sigma^{2}=0.1$. We fix the Hurst parameter of fBM to be 0.45 in all experiments.